# Toward a Deterministic Nucleation Theory for Chirality-Controlled Nanotube Synthesis

### v2

@@
Abstract
The electronic properties of carbon nanotubes are dictated by their chirality, specified by the integer indices $(n,m)$. The absence of a predictive theory directly linking catalysts to the resulting $(n,m)$ has hindered the controlled synthesis of specific chiralities. Here we narrow this gap through a two-step approach. First, we derive a universal vector sum rule that quantifies the relationship between cap structure and chirality, demonstrating that $(n,m)$ is uniquely determined by the vector sum of the six pentagon coordinates within the cap. The complete set of these coordinates, evaluated across rotationally symmetric lattice frames, fully defines the carbon cap structure. Building on this topological foundation, we establish that chirality is encoded deterministically at nucleation, as perturbations to pentagon positions lead to predictable, quantifiable, and kinetically irreversible shifts in $(n,m)$. By examining existing experimental evidence, we further propose that the preferential formation of $(12,6)$ nanotubes arises from a six-fold symmetric cap with epitaxial matching to the $\langle 1,1,1\rangle$ catalyst facet. Our work provides a theoretical framework that redefines the field, shifting the paradigm from stochastic growth kinetics to deterministic nucleation programming and paving the way toward predictable synthesis.
@@ Abstract
Carbon nanotubes (CNTs) hold exceptional promise for semiconductor applications due to their extraordinary electrical conductivity and chirality-dependent band gaps [1, 2]. However, achieving chiral-selective synthesis remains a fundamental challenge in nanotechnology [3, 4, 5]. Chemical vapor deposition with solid catalysts offers a scalable and promising approach toward chirality-specific growth, having produced several CNTs with remarkable enantiomeric purity—including $(12,6)$ at 92% and $(14,4)$ at 99% [6, 7]. Despite these advances, the physical principles governing chiral preference remain poorly understood. Moreover, the absence of a predictive theory connecting catalysts to resulting chiralities has impeded systematic extension of this approach to other chiralities [8, 9, 10, 11, 12, 13].
Carbon nanotube growth proceeds via crystalline elongation from a nucleation cap—a curved graphene dome formed by incorporating six pentagons into a hexagonal lattice [14]. This cap architecture encodes complete chirality information within its topological structure [15, 16]. Developing a predictive theory connecting solid catalysts to the chiral outcomes requires addressing two interconnected challenges: deciphering chirality from cap topology, and elucidating how catalysts direct cap formation during nucleation. The first is a problem of topological interpretation, while the second involves complex physics of nucleation and catalyst-carbon interactions. Resolving these intertwined problems is essential for advancing this important field.
@@
Figure 1: Schematic of a carbon cap structure. Red pentagons mark the positions of the pentagon carbon rings. Top left: the lattice basis vectors. Bottom left: the three topological basis vectors used in our framework.
@@
$\overset{\leftrightarrow}{\mathbf{P}}_{6}=\mathbf{e}_{j}\cdot\mathbf{e}_{k}\cdot\mathbf{e}_{i}^{-1}\cdot\mathbf{e}_{j}^{-1}\cdot\mathbf{e}_{k}^{-1}\cdot\mathbf{e}_{i}.$
@@
In graphene or nanotubes, any topological path connecting crystallographically equivalent atoms ($A\rightarrow A$ or $B\rightarrow B$) belongs to the group $V$. The value of such a path is defined as the element $\alpha\mathbf{p}_{ji}+\beta\mathbf{p}_{jk}$ (with $\alpha,\beta\in\mathbb{Z}$) to which it is homotopic. This leads to a purely topological definition of chirality: for a carbon nanotube of chirality $(n,m)$, a closed path encircling the circumference takes the value $n\mathbf{p}_{ji}+m\mathbf{p}_{jk}$ (see Supplemental Material S1 for detailed description of the topological framework).
@@
Figure 2: Chiral index assignment of a carbon nanotube cap. (Top left) Six rotationally generated vector spaces obtained by successive left rotations. (Bottom) Chiral indices of the cap mapped onto the graphene lattice in corresponding vector spaces. (Top right) Atomic configuration of the cap and the resulting carbon nanotubes, with the cap-nanotube boundary defined by the presence of a stable edge structure during nucleation (see Supplemental Material S2 for details)
@@
Figure 3: (a)-(c) Schematic illustrations of lattice matching between nanotube caps and $\langle 1,1,1\rangle$ crystallographic facets of catalyst nanoparticles. Cap notations $\ldots], [(2,1),(2,1),$(3,0),(2,1),(2,1),(3,0)][(2,1),(2,1),(3,0),(2,1),(2,1),(3,0)], and $(3,0),\ldots]$ correspond to $(12,6)$, $(14,4)$, and $(18,0)$ nanotubes, respectively. (d) Per-carbon-atom formation energy of caps on copper catalyst surface as a function of chiral angle $\theta$, determined by molecular dynamics simulations enumerating 730 distinct configurations corresponding to 18 chiralities. The initial graphene lattice contains 19 hexagons, with pentagon incorporation occurring externally at the edge. Atomic interactions are described by a bond-order potential [18], under NVT ensemble with temperature maintained at 1 K throughout all simulations. The color bar shows the nucleation probabilities derived from a simplified model.
Two distinct pathways—clone growth and solid catalyst–specified growth via CVD—offer routes toward atomic-level control over nanotube synthesis. Clone growth directly programs the cap architecture through molecular-scale design. For instance, chemical opening of $\mathrm{C}_{60}$ molecules yields well-defined cap sequences such as $(1,0),(2,0),\ldots]$ and $(0,0),(1,1),\ldots]$, which template the growth of $(9,0)$ and $(5,5)$ nanotubes, respectively (Fig. 2[19]. Alternatively, molecular templates synthesized through organic chemistry—such as the $\mathrm{C}_{96}\mathrm{H}_{54}$ molecule, which preorganizes the sequence $(2,1),(0,1),\ldots]$—can direct cap assembly to selectively form $(6,6)$ nanotubes [20]. Clone growth provides a clear demonstration of the direct pathway from nucleation to nanotube formation. However, it suffers from slow elongation rates beyond the microscale.
@@
First, we need to establish the concept of deterministic nucleation by addressing a fundamental question: is chirality a statistical outcome of stochastic nucleation processes, or is it deterministically determined by a specific nucleation process that directs growth toward a preferred $(n,m)$? This question is subtle yet critically important. The unconscious assumption of the former has led to an overwhelming focus on post-nucleation explanations of chirality control, particularly through edge-matching mechanisms. This issue relates to the fact that each chirality corresponds to a large number of distinct cap configurations, with this number growing rapidly with diameter. Eq. (6) establishes that any change in pentagon position induces a deterministic and quantifiable shift in chirality, unless compensated by countervailing changes elsewhere in the cap. For instance, shifting one pentagon from $(0,0)$ to $(1,1)$ in a $(5,5)$ nanotube redirects growth to a $(6,6)$ nanotube (Fig. 2). Restoring the original $(5,5)$ chirality would require compensating changes in other pentagon positions summing to (1,1)(-1,-1)—a scenario of low statistical probability in an entropy-dominated regime. This implies that if the nucleation process equally generated all possible caps for subsequent growth, it would also equally generate caps for neighboring chiralities through minor pentagon position shifts. Such behavior is never observed experimentally. Chirality enrichment follows specific patterns that cannot be explained by proximity in chirality space. For example, while $(12,6)$ nanotubes are enriched, $(12,7)$ nanotubes have not been observed enriching in any experimental studies. From a broader perspective, the total number of caps meeting the diameter requirements proximate to $(12,6)$ nanotubes reaches millions, yet only a very small percentage correspond specifically to $(12,6)$. Statistically, it is impossible for a random nucleation process to result in such chirality enrichment. The origin of chirality control must therefore stem from deterministic nucleation where only a few (or one) closely related caps are engineered—as unambiguously evidenced by the precise structural templating observed in clone growth experiments [19, 20].
The remaining challenges are to identify the specific cap architectures responsible for the preferential formation of $(12,6)$ nanotubes and to understand the underlying physical principles that favor such configurations. The vector sum rule provides the foundation to systematically enumerate all possible caps satisfying this chirality condition under catalyst constraints, enabling targeted investigation of their relative stabilities and catalytic matching properties..
carbon caps exhibiting six-fold rotational symmetry. These caps correspond to nanotubes with chiral indices $(6n,6m)$ and are topologically represented by the notation $(n,m),\ldots]$, as exemplified by $(1,0),\ldots]$, $(1,1),\ldots]$, and $(2,1),\ldots]$. Such symmetric caps achieve coherent structural registry with $\langle 1,1,1\rangle$ crystallographic facets of various face-centered catalysts, with all six pentagons positioned along the cap periphery (Fig. 3a). This configuration creates favorable nucleation conditions, enabling epitaxial extension of the cap across the catalyst surface in a graphene-like growth mode [25, 26]. Pentagon incorporation is triggered either when the expanding carbon network encounters the catalyst boundary or when progressive lattice mismatch during epitaxial growth induces large-scale structural strain [27, 28, 29, 30].
@@
The enrichment of $(12,6)$ nanotubes can thus be explained by the combination of optimal nucleation characteristics with favorable post-nucleation growth kinetics—specifically, a diameter of approximately 1.3 nm and a chiral angle of $19.1^{\circ}$ [31, 32, 33].
@@
However, achieving such well-defined growth requires precise incorporation of all six pentagons at $(2,1)$ sublattice sites (or equivalently, $(1,2)$ sites, which yield the mirror-symmetric $(6,12)$ chirality). Any deviation from this ideal configuration—for example, substituting two $(2,1)$ pentagons with $(3,0)$ sites—deterministically shifts the resulting chirality to the semiconducting $(14,4)$ type (Fig. 3b). Consistent with this theoretical framework, the $(14,4)$ nanotube exhibits experimental enrichment under similar synthesis conditions [7]. Moreover, several other structurally related chiralities—including $(18,0)$, $(16,2)$, $(15,3)$, and (9,9)(9,9)—are also enriched under specific synthesis conditions [29].
Yet a deeper question persists: why are $(12,6)$ nanotubes overwhelmingly preferred over other chiralities across different synthesis conditions? Although a rapid post-nucleation elongation rate makes them easier to detect experimentally, this alone cannot explain their abundance across diverse synthesis conditions unless preferential nucleation conditions also exist relative to neighboring caps. Among the 730 distinct cap configurations (satisfying the Isolated Pentagon Rule) closely related to this six-fold symmetry, only 116 configurations yield $(12,6)$ chirality. From a configurational entropy perspective alone, other chiralities should be substantially more probable within this symmetry family. The experimental predominance of $(12,6)$ nanotubes therefore points to decisive energetic advantages and kinetic privilege of the $(2,1),\ldots]$ cap that override entropic factors.
As revealed by in situ transmission electron microscopy [26], during nucleation, edges undergo repeated etching and regrowth, gradually adjusting toward lower energy states. In this stage, low-energy cap configurations gain significant advantage over alternatives. As shown in Fig. 3c, the formation energy of caps registered with crystallographic facets approximately follows a scaled sinusoidal function of the chiral angle, with the $(2,1),\ldots]$ cap exhibiting the largest deviation from this trend. This particular advantage of the $(2,1),\ldots]$ cap arises from two key factors: maximized inter-pentagon distances that reduce strain energy, and closer registry with catalyst surfaces that minimizes interfacial energy. Furthermore, nucleation is not entirely an equilibrium process; the kinetics of pentagon incorporation also play a crucial role. Experimental evidence indicates that incorporating a pentagon by adding a $\text{C}_{3}$ atom sequence to an anti-armchair site is significantly less favorable than adding a $\text{C}_{2}$ atom to a zigzag site [30]. Consequently, the incorporation of a pentagon at a $(3,0)$ site requires either its right neighboring $(1,2)$ site or left neighboring $(2,1)$ site to incorporate a hexagon first, thereby converting the $(3,0)$ site to a zigzag configuration.
To quantify this kinetic barrier, we consider a simplified model where the probability of pentagon incorporation at a zigzag site is $\phi$ ($0<\phi<1$), while the probability of hexagon incorporation at the same site is $1-\phi$. The overall probability of pentagon incorporation at a $(3,0)$ site then becomes $2(1-\phi)\phi$. Because incorporating a hexagon at a zigzag site also requires a $\text{C}_{3}$ atom sequence, which is less favorable than the $\text{C}_{2}$ addition for pentagon incorporation, we have $1-\phi<\phi$. Therefore, the probability $2(1-\phi)\phi$ is lower than the probability $\phi$ for direct pentagon incorporation at $(2,1)$ or $(1,2)$ sites. The total probability to form a cap, ignoring probability fluctuations from other incorporation events, is given by $\Psi=\prod_{t=1}^{6}\phi_{t}$, where $\phi_{t}$ represents the incorporation probability at each of the six pentagon sites. $\Psi$ represents the relative formation probability of a specific cap compared to other configurations. This simple kinetic analysis confirms that the $(2,1),\ldots]$ cap is privileged and favored, while pro-zigzag caps are significantly less favorable to nucleate despite having lower formation energy in some cases, as illustrated in Fig. 3c (assuming $\phi=0.9$ for demonstration purposes). From a stochastic perspective, although chirality is deterministically encoded by each cap, the emergence of a distribution of chiralities during synthesis is statistically inevitable, given that the probability to form unfavored caps is low but non-zero. Of course, an accurate model needs to consider all incorporation pathways and, most importantly, account for how morphological changes of the catalyst surface influence the incorporation probabilities. This critical issue has been a main focus in graphene synthesis toward large-scale pristine lattice, being studied by in situ transmission electron microscopy [30]. In graphene studies, pentagons and heptagons typically emerge as pairs, whereas in nanotubes, the nucleation process involves exclusively pentagons. This remains an uncharted territory; therefore, a concise study of pentagon incorporation dynamics is required [34, 35].
Beyond nucleation control, our topological framework opens pathways for the rational design of carbon-based heterostructures, including intramolecular nanotube junctions and graphene-nanotube hybrids (see Supplemental Material S3). By strategically incorporating pentagonal and heptagonal defects into the hexagonal lattice, diverse target architectures can be systematically constructed. Thus, our theory serves as a versatile and predictive tool for the atomic-scale engineering of complex carbon nanomaterials.

### v3

@@
Toward a Deterministic Nucleation Theory for Chirality-Controlled Nanotube Synthesis
Zhengrong Guo zhengrong_guo@yulinu.edu.cn
[…]
The electronic properties of carbon nanotubes are governed by their chirality, specified by the integer indices $(n,m)$. While chirality-controlled synthesis has achieved notable successes, theoretical understanding remains predominantly focused on post-nucleation growth. Two fundamental obstacles impede deeper insight: the absence of a clear description of nucleation cap topology and its connection to tube chirality, and an incomplete understanding of atomic-level mechanisms governing templated cap formation. Here we address these challenges directly. First, we develop a mathematically rigorous topological framework for carbon networks that provides both a concise definition of cap structures and a quantitative relationship between cap architecture and chirality—the vector sum rule. Second, contrary to conventional perspectives attributing chirality enrichment to edge matching during growth, we demonstrate that chirality is deterministically encoded during nucleation through selective formation of specific cap structures on catalyst surfaces. For the specific case of $(12,6)$ nanotubes, we show that their enrichment arises from a six-fold symmetric cap with epitaxial matching to catalyst facets. Our deterministic nucleation theory not only provides a coherent explanation for chirality enrichment but also elucidates its pattern in chirality space. This work establishes a theoretical framework that redefines the field, shifting the paradigm from stochastic growth kinetics to deterministic nucleation programming and paving the way toward predictable synthesis.
@@
Carbon nanotubes (CNTs) hold exceptional promise for semiconductor applications due to their extraordinary electrical conductivity and chirality-dependent band gaps [1, 2]. However, achieving chiral-selective synthesis remains a fundamental challenge in nanotechnology [3, 4, 5]. Chemical vapor deposition with solid catalysts offers a scalable and promising approach toward chirality-specific growth, having produced several CNTs with remarkable enantiomeric purity—including $(12,6)$ at 92% and $(14,4)$ at 97% [6, 7]. Despite these advances, the physical principles governing chiral preference remain poorly understood. Moreover, the absence of a predictive theory connecting catalysts to resulting chiralities has impeded systematic extension of this approach to other chiralities [8, 9, 10, 11, 12, 13]. Nanotube growth proceeds via crystalline elongation from a nucleation cap—a curved graphene dome formed by incorporating six pentagons into a hexagonal lattice [14]. This cap architecture encodes complete chirality information within its topological structure [15, 16]. Developing a predictive theory connecting solid catalysts to the chiral outcomes requires addressing two interconnected challenges: deciphering chirality from cap topology, and elucidating how catalysts direct cap formation during nucleation. The first is a problem of topological interpretation, while the second involves complex physics of nucleation and catalyst-carbon interactions. Resolving these intertwined problems is essential for advancing this important field.
@@
Figure 1: Schematic of a carbon cap structure. Top left: the lattice basis vectors. Bottom left: the three topological basis vectors used in our framework. Red pentagons mark the positions of the pentagon carbon rings.
@@
${\leftrightarrow}{\mathbf{P}}_{6}=\mathbf{e}_{j}\cdot\mathbf{e}_{k}\cdot\mathbf{e}_{i}^{-1}\cdot\mathbf{e}_{j}^{-1}\cdot\mathbf{e}_{k}^{-1}\cdot\mathbf{e}_{i}.$
@@
In graphene or nanotubes, any topological path connecting crystallographically equivalent atoms ($A\rightarrow A$ or $B\rightarrow B$) belongs to the group $V$. The value of such a path is defined as the element $\alpha\mathbf{p}_{ji}+\beta\mathbf{p}_{jk}$ (with $\alpha,\beta\in\mathbb{Z}$) to which it is homotopic. This leads to a purely topological definition of chirality: for a carbon nanotube of chirality $(n,m)$, a closed path encircling the circumference takes the value $n\mathbf{p}_{ji}+m\mathbf{p}_{jk}$ (see Supplemental Material for detailed description of the topological framework).
@@
Figure 2: Chiral index assignment of a carbon nanotube cap. (Top left) Six rotationally generated vector spaces obtained by successive left rotations. (Bottom) Chiral indices of the cap mapped onto the graphene lattice in corresponding vector spaces. (Top right) Atomic configuration of the cap and the resulting carbon nanotubes.
@@
Two distinct pathways—clone growth and solid catalyst–specified growth via CVD—offer routes toward chirality control over nanotube synthesis. Clone growth directly programs the cap architecture through molecular-scale design. For instance, chemical opening of $\mathrm{C}_{60}$ molecules yields well-defined cap sequences such as $(1,0),(2,0),\ldots]$ and $(0,0),(1,1),\ldots]$, which template the growth of $(9,0)$ and $(5,5)$ nanotubes, respectively (Fig. 2[18]. Alternatively, molecular templates synthesized through organic chemistry—such as the $\mathrm{C}_{96}\mathrm{H}_{54}$ molecule, which preorganizes the sequence $(2,1),(0,1),\ldots]$—can direct cap assembly to selectively form $(6,6)$ nanotubes (Fig. 2[19]. Clone growth provides a clear demonstration of the direct pathway from nucleation to nanotube formation. However, it suffers from slow elongation rates beyond the microscale.
@@
We begin by examining a fundamental dichotomy in chirality-controlled nanotube synthesis: is chirality statistically determined by post-nucleation kinetic processes, or deterministically encoded through selective cap formation? While subtle, this distinction carries profound implications. The prevailing assumption of former has directed extensive attention toward post-nucleation chirality control via edge-matching mechanisms. This perspective, however, faces a critical challenge: each chirality corresponds to numerous distinct cap configurations, a number that grows rapidly with diameter, as does the diversity of possible chiralities. Indeed, the configurational entropy of possible cap structures is effectively infinite. Consequently, any uncontrolled nucleation process would yield the complete spectrum of cap structures and, therefore, all possible nanotube chiralities. Even with diameter constraints, numerous chiralities exhibiting similar favorable post-nucleation growth conditions would be nucleated at same time. Thus, post-nucleation mechanisms alone cannot adequately explain the observed selective chirality enrichment. The origin of chirality control must therefore originate from deterministic nucleation processes where only specific cap structures are selectively formed. The solid catalyst serves as a template that dramatically narrows the vast configuration space to a limited set of specific structures—as unequivocally demonstrated by the precise structural templating observed in clone growth experiments [18, 19].The challenge then becomes identifying the specific cap architectures responsible for the preferential formation of $(12,6)$ nanotubes and understanding the underlying physical principles that favor such configurations. The vector sum rule provides the foundation to systematically enumerate all possible caps satisfying this chirality condition under catalyst constraints, enabling targeted investigation of their relative stabilities and catalytic matching properties.
[…]
@@
The enrichment of $(12,6)$ nanotubes can thus be explained by the combination of favorable nucleation characteristics with optimal post-nucleation growth kinetics—specifically, a diameter of approximately 1.3 nm and a chiral angle of $19.1^{\circ}$ [30, 31, 32].
Figure 3: A. Schematic illustrations of cap formation on $\langle 1,1,1\rangle$ crystallographic facets of catalyst nanoparticles. B. Representative cap structures corresponding to distinct chiralities derived from the $\mathrm{C}_{54}$ intermediate. C. Distribution of $\mathrm{C}_{54}$-derived caps mapped in chirality space. (D) Per-carbon-atom formation energy of caps on a copper catalyst as a function of chiral angle $\theta$, with color scale representing nucleation probabilities.
@@
However, achieving such well-defined growth requires precise incorporation of all six pentagons at $(2,1)$ sublattice sites (or equivalently, $(1,2)$ sites, which yield the mirror-symmetric $(6,12)$ chirality). Any deviation from this ideal configuration direct the grow to alternative chirality. For example, substituting two $(2,1)$ pentagons with $(3,0)$ sites—deterministically shifts the resulting chirality to the semiconducting $(14,4)$ type. Consistent with this theoretical framework, the $(14,4)$ nanotube exhibits experimental enrichment under similar synthesis conditions [7]. Moreover, several other structurally related chiralities—including $(18,0)$, $(16,2)$, $(15,3)$, and (9,9)(9,9)—are also enriched under specific synthesis conditions(Fig. 3B) [24].
A fundamental question persists: why do $(12,6)$ nanotubes exhibit such overwhelming dominance under typical synthesis conditions? Consider a solid catalyst with a diameter of 1.4 nm—a common size in nanotube growth—whose surface possesses a facet of 19 atoms arranged in a honeycomb lattice. This specific facet template dictates the exclusive formation of the $\mathrm{C}_{54}$ carbon cluster. The periphery of this cluster presents 18 active sites for pentagon incorporation, generating 730 possible cap configurations ($N_{\text{config}}$) that satisfy the isolated pentagon rule (Fig. 3C). These correspond to 70 structurally distinct caps ($N_{\text{cap}}$) , which in turn seed 10 different nanotube chiralities. If pentagon incorporation were a random process, the resulting chirality distribution would mirror the statistical weights of these configurations—a prediction starkly at odds with experimental evidence. The pronounced enrichment of $(12,6)$ nanotubes from this vast configurational space unequivocally demonstrates that decisive energetic and kinetic factors override entropic considerations.
As revealed by in situ transmission electron microscopy [26], nanotube edges undergo repeated etching and regrowth during nucleation, progressively evolving toward lower energy states. At this stage, low-energy cap configurations gain a significant advantage over alternatives. As shown in Fig. 3C, the $(2,1),\ldots]$ cap exhibits the lowest formation energy among all 730 possible configurations. This distinctive advantage stems from two key factors: maximized inter‑pentagon distances that reduce strain energy, and improved registry with the catalyst surface that minimizes interfacial energy.
However, whether this nucleation process reaches thermodynamic equilibrium remains heavily debated. Previous studies have indeed identified kinetic barriers in the growth of some chiral tubes [15]. Consequently, confirming the absence of kinetic barriers in the nucleation pathway for the $(2,1),\ldots]$ cap is essential. Experimental evidence indicates that the addition of a $\mathrm{C}_{3}$ sequence to an active site is significantly less favorable than the incorporation of a $\mathrm{C}_{2}$ unit. Notably, from a purely kinetic standpoint, pentagon incorporation is generally favored over hexagon formation in most scenarios [29].
The first application of kinetic theory to nanotube growth rate analysis was pioneered by Yakobson et al. [30], who modeled growth rates as proportional to the number of active sites available for $\mathrm{C}_{2}$ addition. We extend this framework by defining the probability of a $\mathrm{C}_{2}$ addition event as $\phi(t)=1-\mathrm{e}^{-\lambda t}$, where $\lambda$ is an experimentally determined rate constant and $t$ is time. Cap formation is considered complete when a new edge is generated on the existing $\mathrm{C}_{54}$ intermediate. To simplify the model, we consider only $\mathrm{C}_{2}$ additions at active sites and treat $\lambda$ as constant for all events. A fundamental constraint is that hexagon incorporation becomes feasible only when an adjacent site has already incorporated a pentagon or hexagon. This leads to an Erlang probability density for completing an $N$-subsequence incorporation at time $t$, described by $\Theta(t,N)=\lambda^{N}t^{N-1}\mathrm{e}^{-\lambda t}/\Gamma(N)$. Each local incorporation event culminates in a hexagon at the $(3,0)$ position, requiring pre-incorporation of neighboring sublattices from both directions. The probability density for this sub-event at time $t$ is given by
$\Psi(t,\alpha,\beta)=\int_{0}^{t}\Theta(\tau,\alpha)\Theta(\tau,\beta)\lambda\mathrm{e}^{-\lambda(t-\tau)}d\tau,$
where $\alpha$ and $\beta$ represent the number of incorporations from each respective direction.
In the simplest case of the $[(3,0),(3,0),(3,0),(3,0),(3,0),(2,1)]$ cap—corresponding to the $(17,1)$ tube—incorporation initiates at the $(2,1)$ sublattice pentagon site and propagates sequentially around the circumference until reaching the final $(3,0)$ sublattice. The cumulative probability distribution is given by $F_{(17,1)}(t)=\int_{0}^{t}\Psi(\tau,16,1)\,d\tau$.
For the symmetric $(2,1),\ldots]$ cap associated with the $(12,6)$ tube, edge expansion begins independently at all six $(2,1)$ sites, culminating in hexagon formation at each $(3,0)$ position. The cumulative formation probability by time $t$ follows as
$F_{(12,6)}(t)=\left[\int_{0}^{t}\Psi(\tau,2,1)\,d\tau\right]^{6},$
where the integral reflects the cumulative probability for each individual growth arm.
In the special case of the $(3,0),\ldots]$ cap corresponding to the $(18,0)$ tube, a $\mathrm{C}_{3}$ addition is required to form the initial hexagon or pentagon, a necessary step for propagation. We therefore treat its incorporation probability as effectively negligible.
Our kinetic model reveals a pronounced advantage of the $(2,1),\ldots]$ cap structure over most alternative configurations (Fig. 3D, parameters: $\lambda=10^{3}$, $t=10^{-2}$, corresponding to a typical nucleation timescale of 10 ms). As nanotube growth is governed by both thermodynamic stability and kinetic accessibility, we demonstrate conclusively that the dominance of the $(2,1),\ldots]$ cap drives the preferential formation of $(12,6)$ nanotubes.
Within this framework, a minority of cap structures are predicted to have marginally higher formation probabilities than the symmetric $(2,1),\ldots]$ cap. A representative case is the $(2,1),(1,2),\ldots]$ cap of the armchair $(9,9)$ tube. This apparent overestimation stems from pathways involving hexagon incorporation between neighboring pentagons—a event that necessitates a substantial structural reorganization and is thus expected to be energetically costly, resulting in a reduced $\lambda$ value. A predictive kinetic model must therefore integrate the complete ensemble of structural variations while accounting for the influence of dynamic catalyst morphology on incorporation probabilities. The pivotal role of catalyst morphology has been extensively demonstrated in the synthesis of pristine graphene through in situ transmission electron microscopy studies [27]. While pentagon-heptagon pairs represent the dominant defects in graphene, nanotube nucleation is governed exclusively by pentagon formation [33, 34]. The dynamics of pentagon incorporation during nucleation remain largely unexplored and warrant detailed investigation.